OpenStudy (anonymous):

7x + 2y =5 ????????????

OpenStudy (fanduekisses):

x= 5/7 y=5/2

OpenStudy (anonymous):

really?ty

OpenStudy (fanduekisses):

yes you just solve the x and y separately

OpenStudy (anonymous):

k ty

OpenStudy (fanduekisses):

for example 7x + 0 = 5 7x = 5 /7 /7 x= 5/7

OpenStudy (fanduekisses):

0 + 2y = 5 2y = 5 /2 /2 y = 5/2

OpenStudy (anonymous):

so no x=5/7y just x=5/7

OpenStudy (anonymous):

now im confused again

OpenStudy (fanduekisses):

The answer is x= 5/7 y= 5/2 you just solve for x and y

OpenStudy (anonymous):

ok

OpenStudy (fanduekisses):

divide 5 by 7 and 5 by 2